OpenStudy (jacobciezki):

help plz :(Evaluate. 2 – 4 • 3 + 6 • 3 A. –28 B. 12 C. 8 D. 0

OpenStudy (ilovebmth1234):

C.8

OpenStudy (djbreezy):

Its 8

OpenStudy (jacobciezki):

how you get that

OpenStudy (djbreezy):

www.mathway.com

OpenStudy (ilovebmth1234):

(4*3)= 12 (6*3)= 18 so 2- 12 +18= 8

OpenStudy (djbreezy):

just type ur problem it give u the answer

OpenStudy (jacobciezki):

ok

OpenStudy (ilovebmth1234):

right but you cant always use a website for math

OpenStudy (jacobciezki):

this is so hard lol